# Reproducibility and Artifact Consistency of the SIGIR 2022 Recommender Systems Papers Based on Message Passing

### v1

@@ 3.4. Learning to Denoise Unreliable Interactions for Graph Collaborative Filtering
RGCF is evaluated on three datasets: Yelp2018, Amazon-Book and Movielens 1M. Both Yelp2018 and Amazon-Book are preprocessed with a 15-cores selection, while for Movielens 1M only interactions whose rating is at least 4 are retained and associated to an implicit rating of 1. The data splitting is a random holdout of interactions sampled globally, 80% training, 10% validation and 10% test.
@@ 3.5. INMO: A Model-Agnostic and Scalable Module for Inductive Collaborative Filtering
The GitHub repository [^24] https://github.com/WuYunfan/igcn_cf contains the implementation, the training-test data split, and, uniquely among the algorithms analyzed in this study, it also includes the implementation of the baseline models along with their own hyperparameter optimization code. The provided material is fully consistent with what is described in the paper. [^25] We note that the paper uses the BPR loss and defines it by using the log sigmoid function:$-ln(\sigma(x))$, while the implementation instead uses the $\mbox{softplus}(-x)=ln(1+e^{-x})$ function. This occurs in a few of the papers we analyze and we wish to point out that the two formulations are mathematically identical: $-ln(\sigma(x))=-ln\left(1/(1+e^{-x})\right)=ln(1+e^{-x})=\mbox{softplus}(-x)$. As a minor note, the paper reports a number of interactions for all datasets that is approximately 15% higher than that of the provided data. The number of epochs is correctly selected using early-stopping, where training stops if the NDCG@20 on the validation data does not improve for 50 successive epochs.
@@ 3.9. Knowledge Graph Contrastive Learning for Recommendation
3.10. Comparison of Between the SIGIR 2022 Analyzed Methods
To provide a more complete picture of model effectiveness and to separate our assessment from potential confounding factors related to hyperparameter optimization, which, as we have pointed out, is generally extremely limited, we select two datasets and conduct an independent hyperparameter optimization. It is important to note that this type of comparison comes with a significant computational cost, as all the examined methods require iterative training. The experiments we report required approximately six months of GPU time, making this type of evaluation infeasible for many academic or independent researchers.
@@ Results
The results for the Amazon-Book dataset, reported in Table 12, once again show that none of the GNN methods outperform our selection of simple baselines. However, we observe that INMO matches SLIM in terms of Recall and that, for some methods (INMO, KGCL, SimGCL, and LightGCN), the effectiveness gap is much smaller compared to the previous results we obtained in our reproducibility analysis, see Section 3.5, 3.9, 3.3 and 3.1. On the other hand, both HCCF and GTN prove ineffective. While HCCF completed the allotted 50 hyperparameter trials, GTN was limited to only 15 due to its long training time which may partly explain its poor effectiveness. Again, as we previously observed in Section 3.2, GDE exhibits numerical instabilities that prevented it from training properly. Note that slight differences in preprocessing and data splitting make it impossible to directly compare the absolute values of effectiveness metrics across many of the experiments we reported in the previous sections. For example, in the analysis of SimGCL (see Section 3.3), its results with the original hyperparameters showed an NDCG@20 of 0.1047, whereas SLIM achieved 0.1816, a substantial difference. However, in this comparison experiment on Amzon-Book, the two methods are within 2%. A similar trend is observed for INMO, which previously had a result of 0.0934 compared to SLIM’s 0.1451, a 55% difference, while in this experiment the results differ by 7%; KGCL, which reached 0.0794 compared to SLIM’s 0.1031, a 30% difference, while in this experiment results differ by 11%; and LightGCN, which originally reached 0.0862 compared to SLIM’s 0.1838, a whopping 113% difference, while here the results differ by 11%. Given that the data split used here does not appear to exhibit anomalies, we argue that a major reason for this large difference is the ineffectiveness of the original hyperparameter tuning. Since the weak baselines used for comparison did not present a sufficient challenge, there was little incentive to properly fine-tune the methods. This further reinforces the observation that the experimental practices adopted in the community are often ineffective and unreliable, and allow researchers to support many contradicting conclusions depending on minor changes in the experimental protocol. As discussed by Shehzad and Jannach (2023), ”Everyone’s a Winner”, any method can appear competitive if the analysis is conducted carelessly, and current peer review practices do not seem capable of detecting such cases.
@@ 4.1. Artifacts
RGCF Amazon-Book, Movielens 1M, Yelp2018 15383838With the exception of Movielens 1M, where no k-cores is applied. global 80% - 10% - 10% No -
@@ 4.5. Impact on Follow-up Research
Our goal for this section is to assess to what extend irreproducible or weak methods from SIGIR 2022 have been used as baselines in SIGIR 2023. Table 16 lists, for each of the eleven SIGIR 2023 papers we identified, the SIGIR 2022 papers they used as baselines. It is clear that LightGCN is the most influential model, included as baseline in eight out of eleven papers. The second most frequently used baseline is SimGCL, included in four papers, followed by HCCF, which is used in three. In all of these cases, the methods are presented as representing the state-of-the-art, which, as we have observed in this study is not true with very limited exceptions. In our independent optimization, see Section 3.10, HCCF and LightGCN were consistently below the baseline models and SimGCL exhibited competitive results only on Yelp2018. Unfortunately, it is challenging to compare and cross-check the results for the baselines with those reported in the papers that proposed them. For example, one might be interested to assess the consistency in the effectiveness of LightGCN, SimGCL, and HCCF, the three most-used baselines. In their original papers, LightGCN and SimGCL employ splits of 72%-8%-20% on Gowalla, Amazon-Book, and Yelp2018. HCCF, on the other hand, is evaluated on Yelp, Amazon-Book, and MovieLens-10M with a 70%-10%-20% split. When comparing this with the evaluation procedure used in the SIGIR 2023 papers we find that none adopts a data split consistent with those used in the original baseline papers. Furthermore, one paper (Wang et al., 2023) uses a 5-fold cross-validation instead of a single hold-out training-test set, while two papers Yang et al. (2023); He et al. (2023) repeat the experiments 5 times on the same test set. Although these are good practices which allow to measure the variance of the results, they further complicate direct comparisons between papers especially when the variance is not reported. On one hand it is somewhat beneficial that the erroneous preprocessing and splitting steps used by LightGCN and SimGCL do not appear to be adopted in the SIGIR 2023 papers, except possibly (Choi et al., 2023). [^52] It should be noted that this cannot be confirmed without directly verifying the content of the data splits. However, this comes at the significant cost of losing any hope of comparing results across methods and introduces the continuous risk of new mistakes emerging and propagating before they are identified. Indeed, the outcome of this part of the analysis is that it is not possible to compare the results of the SIGIR 2022 papers we analyzed with those reported as baselines in the SIGIR 2023 papers, which we find worrying.
@@
A further step we can take is to examine the results more qualitatively to assess whether the method produces similar results under similar data splitting and preprocessing. However, we find this to be highly challenging. Based on the descriptions provided in the papers, we can identify only a handful of cases where such a comparison is possible. While additional comparable protocols may exist, the lack of sufficiently detailed descriptions in the papers means that discovering them would require manually analyzing the provided data. This type of investigative work is an unreasonable expectation for a researcher simply interested in understanding or building upon a paper. The first case is Wei et al. (2023), which applies a similar data split to SimGCL (Yu et al., 2022), holding out 20% of the data for testing but using different partitions for training and validation. The only dataset the two papers have in common is Yelp2018, and the only shared baseline is SimGCL itself. We observe a notable discrepancy in the results for SimGCL, with its NDCG@20 equal to 0.0601 in the original paper but 0.0795 in (Wei et al., 2023). Due to the limited details provided on hyperparameter optimization and the use of fixed values for some parameters, we are unable to precisely determine whether this difference stems from variations in the data splits or differences in the optimization process. A more successful comparison can be made with Yang et al. (2023), which reports the results of SimGCL and LightGCN on the Douban-Book dataset. Their NDCG@20 values are nearly identical up to the second decimal place. For example, SimGCL achieves 0.1583 in the original paper and 0.1540 in (Yang et al., 2023), while LightGCN reaches 0.1272 in (Yu et al., 2022) and 0.1278 in (Yang et al., 2023). The results for another baseline, MultVAE, are similarly close, with 0.1103 in SimGCL and 0.1155 in (Yang et al., 2023). However, all of these values remain significantly lower than the 0.1694 we obtained for MultVAE in our experiments, which would also slightly outperform the 0.1638 reported for the newly proposed algorithm. The last paper with a similar data split is Li et al. (2023), but once again, we observe effectiveness measurements with substantially different absolute values. For example, the original NDCG@20 for HCCF on the Yelp2018 dataset was 0.0510, while (Li et al., 2023) reports a lower value of 0.0391. A similar discrepancy is found for LightGCN, where the original result on the (erroneous) Yelp2018 split was 0.0530, whereas in (Li et al., 2023), it is 0.0373.
@@
Despite being a research field that heavily relies on experimental findings, these findings seem to often provide little guidance for navigating the vast body of available literature. One of the great successes of the scientific method has been moving beyond an authoritative conception of science, where arguments were deemed valid based on the prestige of the scholar presenting them. However, in a landscape where reliable points of reference are scarce, results are inconsistent and weak models are frequently published, even at top-tier venues, there is a growing risk that readers will be driven to mainly rely on the reputation of researchers or labs as a proxy to assess the correctness of the paper. This is the opposite of what the scientific process is meant to achieve.
@@ 5. Conclusion
The findings of this study highlight significant issues related to artifact consistency and the propagation of poor practices in subsequent publications. Specifically, many of the available training-test data artifacts were clearly the result of an erroneous splitting procedures and were, at the very least, inconsistent with the descriptions provided in the corresponding papers. These data splits, sometimes reused from previous papers, lead to anomalous and questionable evaluation results. When the data split does not appear anomalous, it is typically a unique split created by the authors by applying a specific combination of preprocessing and training-validation-test split percentages, often not clearly described, making comparisons between papers impossible. Indeed, the phenomenon is so prevalent that it is not possible to compare results even between the papers published at SIGIR 2022 and 2023, which we find worrying considering how much of the research literature is based on empirical results. Our analysis also uncovered several bad practices, ranging from arbitrarily setting the values of critical hyperparameters to selecting the number of epochs based on the test data causing information leakage.

### v2

@@ 3.4. Learning to Denoise Unreliable Interactions for Graph Collaborative Filtering
RGCF is evaluated on three datasets: Yelp2018, Amazon-Book and MovieLens 1M. Both Yelp2018 and Amazon-Book are preprocessed with a 15-core selection, while for MovieLens 1M only interactions whose rating is at least 4 are retained and associated to an implicit rating of 1. The data splitting is a random holdout of interactions sampled globally, 80% training, 10% validation and 10% test.
@@ 3.5. INMO: A Model-Agnostic and Scalable Module for Inductive Collaborative Filtering
[…]
@@ 3.9. Knowledge Graph Contrastive Learning for Recommendation
3.10. Comparison Between the SIGIR 2022 Analyzed Methods
To provide a more complete picture of the effectiveness of the GNN models we analyze and to separate this assessment from potential confounding factors related to hyperparameter optimization, which, as we have pointed out, is generally extremely limited, we select two datasets and conduct an independent hyperparameter optimization. It is important to note that this type of comparison comes with a significant computational cost, as all the examined methods require iterative training. The experiments we report in this section required approximately six months of GPU time, making this type of evaluation infeasible for many academic or independent researchers.
@@ Results
The results for the Amazon-Book dataset, reported in Table 12, once again show that none of the GNN methods outperform our selection of simple baselines. However, we observe that INMO matches SLIM in terms of Recall and that, for some methods (INMO, KGCL, SimGCL, and LightGCN), the effectiveness gap is much smaller compared to the previous results we obtained in our reproducibility analysis, see Section 3.1, 3.3, 3.5, and 3.9. On the other hand, both HCCF and GTN prove ineffective. While HCCF completed the allotted 50 hyperparameter trials, GTN was limited to only 15 due to its long training time which may partly explain its poor effectiveness. Again, as we previously observed in Section 3.2, GDE exhibits numerical instabilities that prevented it from training properly. Note that slight differences in preprocessing and data splitting make it impossible to directly compare the absolute values of effectiveness metrics across many of the experiments we reported in the previous sections. For example, in the analysis of SimGCL (see Section 3.3), its results with the original hyperparameters showed an NDCG@20 of 0.1047, whereas SLIM achieved 0.1816, a substantial difference. However, in this comparison experiment on Amzon-Book, the two methods are within 2%. A similar trend is observed for INMO, which previously had a result of 0.0934 compared to SLIM’s 0.1451, a 55% difference, while in this experiment the results differ by 7%; KGCL, which reached 0.0794 compared to SLIM’s 0.1031, a 30% difference, while in this experiment results differ by 11%; and LightGCN, which originally reached 0.0862 compared to SLIM’s 0.1838, a whopping 113% difference, while here the results differ by 11%. Given that the data split used here does not appear to exhibit anomalies, we argue that a major reason for this large difference is the ineffectiveness of the original hyperparameter tuning. Since the weak baselines used for comparison did not present a sufficient challenge, there was little incentive to properly fine-tune the methods. This further reinforces the observation that the experimental practices adopted in the community are often ineffective and unreliable, and allow researchers to support many contradicting conclusions depending on minor changes in the experimental protocol. As discussed by Shehzad and Jannach (2023), ”Everyone’s a Winner”, any method can appear competitive if the analysis is conducted carelessly, and current peer review practices do not seem capable of detecting such cases.
@@ 4.1. Artifacts
RGCF Amazon-Book, MovieLens 1M, Yelp2018 15383838With the exception of MovieLens 1M, where no k-core is applied. global 80% - 10% - 10% No -
@@ 4.5. Impact on Follow-up Research
Our goal for this section is to assess to what extent irreproducible or weak methods from SIGIR 2022 have been used as baselines in SIGIR 2023. Table 16 lists, for each of the eleven SIGIR 2023 papers we identified, the SIGIR 2022 papers they used as baselines. It is clear that LightGCN is the most influential model, included as baseline in eight out of eleven papers. The second most frequently used baseline is SimGCL, included in four papers, followed by HCCF, which is used in three. In all of these cases, the methods are presented as representing the state-of-the-art, which, as we have observed in this study is not true with very limited exceptions. In our independent optimization, see Section 3.10, HCCF and LightGCN were consistently below the baseline models and SimGCL exhibited competitive results only on Yelp2018. Unfortunately, it is challenging to compare and cross-check the results for the baselines with those reported in the papers that proposed them. For example, one might be interested to assess the consistency in the effectiveness of LightGCN, SimGCL, and HCCF, the three most-used baselines. In their original papers, LightGCN and SimGCL employ splits of 72%-8%-20% on Gowalla, Amazon-Book, and Yelp2018. HCCF, on the other hand, is evaluated on Yelp, Amazon-Book, and MovieLens 10M with a 70%-10%-20% split. When comparing this with the evaluation procedure used in the SIGIR 2023 papers we find that none adopts a data split consistent with those used in the original baseline papers. Furthermore, one paper (Wang et al., 2023) uses a 5-fold cross-validation instead of a single hold-out training-test set, while two papers Yang et al. (2023); He et al. (2023) repeat the experiments 5 times on the same test set. Although these are good practices which allow to measure the variance of the results, they further complicate direct comparisons between papers especially when the variance is not reported. On one hand it is somewhat beneficial that the erroneous preprocessing and splitting steps used by LightGCN and SimGCL do not appear to be adopted in the SIGIR 2023 papers, except possibly (Choi et al., 2023). [^52] It should be noted that this cannot be confirmed without directly verifying the content of the data splits. However, this comes at the significant cost of losing any hope of comparing results across methods and introduces the continuous risk of new mistakes emerging and propagating before they are identified. Indeed, the outcome of this part of the analysis is that it is not possible to compare the results of the SIGIR 2022 papers we analyzed with those reported as baselines in the SIGIR 2023 papers, which we find worrying.
@@
A further step we can take is to examine the results more qualitatively to assess whether the method produces similar results under similar data splitting and preprocessing. However, we find this to be highly challenging. Based on the descriptions provided in the papers, we can identify only a handful of cases where such a comparison is possible. While additional comparable protocols may exist, the lack of sufficiently detailed descriptions in the papers means that discovering them would require manually analyzing the provided data. This type of investigative work is an unreasonable expectation for a researcher simply interested in understanding or building upon a paper. The first case is Wei et al. (2023), which applies a similar data split to SimGCL (Yu et al., 2022), holding out 20% of the data for testing but using different partitions for training and validation. The only dataset the two papers have in common is Yelp2018, and the only shared baseline is SimGCL itself. We observe a notable discrepancy in the results for SimGCL, with its NDCG@20 equal to 0.0601 in the original paper but 0.0795 in (Wei et al., 2023). Due to the limited details provided on hyperparameter optimization and the use of fixed values for some hyperparameters, we are unable to precisely determine whether this difference stems from variations in the data splits or differences in the optimization process. A more successful comparison can be made with Yang et al. (2023), which reports the results of SimGCL and LightGCN on the Douban-Book dataset. Their NDCG@20 values are nearly identical up to the second decimal place. For example, SimGCL achieves 0.1583 in the original paper and 0.1540 in (Yang et al., 2023), while LightGCN reaches 0.1272 in (Yu et al., 2022) and 0.1278 in (Yang et al., 2023). The results for another baseline, MultVAE, are similarly close, with 0.1103 in SimGCL and 0.1155 in (Yang et al., 2023). However, all of these values remain significantly lower than the 0.1694 we obtained for MultVAE in our experiments, which would also slightly outperform the 0.1638 reported for the newly proposed algorithm. The last paper with a similar data split is Li et al. (2023), but once again, we observe effectiveness measurements with substantially different absolute values. For example, the original NDCG@20 for HCCF on the Yelp2018 dataset was 0.0510, while (Li et al., 2023) reports a lower value of 0.0391. A similar discrepancy is found for LightGCN, where the original result on the (erroneous) Yelp2018 split was 0.0530, whereas in (Li et al., 2023), it is 0.0373.
@@
Despite being a research field that heavily relies on experimental findings, the growing body of research on reproducibility shows that these published findings often do not reflect the true effectiveness of a method and therefore provide little guidance for navigating the vast body of available literature. One of the great successes of the scientific method has been moving beyond an authoritative conception of science, where arguments were deemed valid based on the prestige of the scholar presenting them. However, in a landscape where reliable points of reference are scarce, results are inconsistent and weak models are frequently published, even at top-tier venues, there is a growing risk that readers will be driven to mainly rely on the reputation of researchers or labs as a proxy to assess the correctness of the paper. This is the opposite of what the scientific process is meant to achieve.
@@ 5. Conclusion
The findings of this study highlight significant issues related to artifact consistency and the propagation of poor practices in subsequent publications. Specifically, many of the available training-test data artifacts were clearly the result of an erroneous splitting procedure and were, at the very least, inconsistent with the descriptions provided in the corresponding papers. These data splits, sometimes reused from previous papers, lead to anomalous and questionable evaluation results. When the data split does not appear anomalous, it is typically a unique split created by the authors by applying a specific combination of preprocessing and training-validation-test split percentages, often not clearly described, making comparisons between papers impossible. Indeed, the phenomenon is so prevalent that it is not possible to compare results even between the papers published at SIGIR 2022 and 2023, which we find worrying considering how much of the research literature is based on empirical results. Our analysis also uncovered several bad practices, ranging from arbitrarily setting the values of critical hyperparameters to selecting the number of epochs based on the test data causing information leakage.